A ticket booth sold 371 tickets and collected $1626.50 in ticket sales. Adult tickets are $5.50 and child tickets are $2.50. How many tickets of each type were sold?
Answer by mananth(16946)

x= number adult tickets
y= number child tickets

1 x + 1 y = 371 .............1
Total value
5.50 x + 2.50 y = 1,626.50 .............2
Eliminate y
multiply (1)by -2.50
Multiply (2) by 1.00
-2.50 x -2.50 y = -927.50
5.50 x + 2.50 y = 1,626.50
Add the two equations
3.00 x = 699.00
/ 3.00
x = 233.00
plug value of x in (1)
1.00 x + 1.00 y = 371.00
233.00 + y = 371.00
y = 371.00 -233
y = 138.00
y = 138.00
x= 233 adult tickets
y= 138 child tickets
m.ananth@hotmail.ca